AutoSave 0 H 2 File Home Insert X cut X Comments U . Exam 2 (Blaschke, Alyssa) (1) - Word Search AO - Draw Design Layout RefeCan you explain how to solve these questions please?
Answer #1

equivalent resistance

R = 35 + 25* ( 45 + 55) / ( 45 + 55 + 25)

R = 55 ohm

=====

current through circuit

i = V/R = 180 / 55 = 3.273 A

=====

v2 = v - v1 = v - i R1 =180 - 3.273*35

v2 = 65.45 V

=====

current through R2

i2 = v2/R2

i2 = 65.45 / 25 = 2.618 A

=======

Comment before rate in case any doubt, will reply for sure.. goodluck